# On the Curvature of Regge Metrics

### v1

@@ 2.5 Removing Frame-Dependency
While useful for analysis, (1) is inconvenient for many applications because it requires knowledge of a specific (generalized) compatible frame, which are not easy to construct or represent. The purpose of this subsection is to remove this dependency by expressing the distributional curvature in terms of endomorphism-valued forms. The cost is that the test forms become discontinuous and metric-dependent.
First, we can define a symmetric nondegenerate bilinear form on the vector space $\mathrm{End}(T_{x}M)=T^{*}_{x}M\otimes T_{x}M$ by setting
$\left\langle\alpha\otimes v,\beta\otimes w\right\rangle=\alpha(w)\beta(v)$ (17)
for $\alpha,\beta\in T_{x}^{*}(M)$, $v,w\in T_{x}M$ and extending multilinearly. Expressed in a basis $\{F_{i}\}_{i=1}^{n}$ and its dual basis $\{F^{i}\}_{i=1}^{n}$, this would be $\left\langle A^{i}_{j}F^{j}\otimes F_{i},B^{k}_{l}F^{l}\otimes F_{k}\right\rangle=A^{i}_{j}B^{j}_{i}$. Therefore, picking any basis for $T_{x}M$ gives a linear isometry between $\mathfrak{gl}(n)$ and $\mathrm{End}(T_{x}M)$, by setting $F^{j}(\hat{\phi}^{T}(F_{i})):=\phi^{j}_{i}$. The left-hand side of this equation could be more succinctly written $([\hat{\phi}]_{F})^{j}_{i}$, so $\hat{\phi}]_{F}$ is the matrix representation of the linear map $\hat{\phi}:T_{x}M\to T_{x}M$ in the basis $F$. In particular, for $y\in B_{T}$, the orthonormal basis $F^{T}(y)$ gives a special isometry $\Psi_{y}^{T}:\mathfrak{gl}(n)\to\mathrm{End}(T_{\Phi^{T}(y)}M)$.
This pointwise isometry in turn defines a map $\Psi_{f}:\bigsqcup_{T\subseteq M}C^{0}\Omega^{n-2}(\mathring{T};\mathfrak{gl}(n))\to\bigsqcup_{T\subset M}C^{0}\Omega^{n-2}(\mathring{T};\mathrm{End}(TM))$, by setting $\Psi_{f}(\phi)|_{x}$ equal to $\Psi_{{\Phi^{T}}^{-1}(x)}^{T}(\phi|_{x})$ if $x$ is in $\mathring{T}$. Clearly, because $\Phi_{y}^{T}$ is invertible for all $y\in\mathring{B_{T}}$ and $\Phi^{T}$ is a diffeomorphism when restricted to $\mathring{B_{T}}$, $\Psi_{f}$ is a bijection, with inverse given by $\Psi_{f}^{-1}(\hat{\phi})|_{x}:=(\Psi_{{\Phi^{T}}^{-1}(x)}^{T})^{-1}(\hat{\phi}|_{x})$.
@@
Each compatible test form (see Definition 2) $\phi$ corresponds to a set of $\mathrm{End}(TM)$-valued forms $\hat{\phi}^{T}\in C^{2}\Omega^{n-2}(\mathring{T};\mathrm{End}(TM))$, which each vanish when pulled back to $\partial M$, defined by $\hat{\phi}|_{x}=\Psi_{{\Phi^{T}}^{-1}(x)}^{T}(\phi|_{x})$ for each codimension-0 polytope $T$. While these forms would usually not fit together into a continuous or even a piecewise-continuous form on $M$, they do have some continuity properties. Since $F^{T}$ is continuous on $B_{T}$ and ${\Phi^{T}}^{*}\phi$ has a unique Lipschitz extension $\tilde{\phi}$ on $B_{T}$, there exists a unique Lipschitz extension of $^{\Phi^{T}}^{*}\hat{\phi}$ to $B_{T}$ by setting ${\Phi^{T}}^{*}\hat{\phi}^{T}|_{y}=\Psi_{y}^{T}(\tilde{\phi}|_{y})$. Note that for codimension-1 faces $e\subset T$, this means the pullback of $\hat{\phi}^{T}$ to $\mathring{e}$ is well-defined, as $i_{\mathring{e}}^{*}\hat{\phi}^{T}={\Phi^{T}}^{-*}i_{{\Phi^{T}}^{-1}(\mathring{e})}^{*}\hat{\phi}^{T}$.
By abuse of notation, we will also use $\hat{\phi}^{T}$ to refer to this extension of $^{\Phi^{T}}^{*}\hat{\phi}$. The difference will be indicated by whether the form is evaluated “on TT” or “on $B_{T}$”.
By (18), at points $x\in e=T\cap T^{\prime}$, we have on $T$ that $i_{e}^{*}\mathrm{Ad}((\mu_{e}^{T})^{-1})([\hat{\phi}^{T}]_{E_{e}^{T}})=i_{e}^{*}\phi=i_{e}^{*}\mathrm{Ad}((\mu_{e}^{T^{\prime}})^{-1})([\hat{\phi}^{T}]_{E^{T^{\prime}}_{e}})$, so
$i_{e}^{*}\hat{\phi}^{T}]_{E_{e}^{T}}=\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{e}^{*}\hat{\phi}^{T^{\prime}}]_{E_{e}^{T^{\prime}}}\right).$ (19)
Similarly, on $B_{T}$, the form $i_{\mathring{p}}^{*}[\hat{\phi}^{T}]_{F^{T}}=i_{\mathring{p}}^{*}\phi$ is well-defined, in the sense that ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}$ does not depend on $s$ (the definition of $\chi_{s,p}^{T}$ is part of definition 2). Below we will use the shorthand $F(s,x)=F^{T}\circ\psi_{p}^{T}(s,x)$.
@@
${\chi_{s,p}^{T}}^{*}[\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}\left(\exp\left(s~r_{p,e}^{T}(x)w^{n-1}_{n}\right)A_{p,e}^{T}\mu_{e}^{T}\right)\left({\chi_{s,p}^{T}}^{*}[\hat{\phi}^{T}]_{F(s,x)}\right)$
@@
Since ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}={\chi_{0,p}^{T}}^{*}\tilde{\phi}^{T}$ is one of the defining characteristics of a compatible $\mathfrak{gl}(n)$-valued test form and ${\chi_{s,p}^{T}}^{*}[\hat{\phi}^{T}]_{F(s,x)}={\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}$, we can thus derive the following equation:
${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}\left(\exp\left(s~r_{p,e}^{T}(x)w^{n-1}_{n}\right)\right)\left([{\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}\right).$ (20)
We will call a set of $\mathrm{End}(TM)$-valued forms which can be obtained this way compatible.
Definition 3.
A compactly supported $\mathrm{End}(TM)$-valued (n2)(n-2)-form $\hat{\phi}$ which is $C^{2}$ on the interior of each codimension-0 polytope $T\subseteq M$ will be called a compatible test form for the compatible frame $f$ if for each $T$, there exists a unique Lipschitz continuous extension of $^{\Phi^{T}}^{*}\hat{\phi}$ to $B_{T}$, equations (19) and (20) are satisfied, and $i_{{\Phi^{T}}^{-1}(\mathring{\Delta_{d}})}^{*}{\Phi^{T}}^{*}\hat{\phi}=0$ if $\Delta_{d}\subset\partial M$. The space of compatible $\mathrm{End}(TM)$-valued test forms for the compatible frame $f$ will be called $\mathcal{A}(f)$.
@@
Just as for compatible test forms, the most important class of compatible $\mathrm{End}(TM)$-valued test forms is given by the forms $\phi^{k}_{lij}(f_{k}\otimes f^{*}\theta^{l})\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$, where the coefficients $\phi^{k}_{lij}$ are continuous compactly supported functions on $M$ which are $C^{2}$ on the interior of each codimension-2 polytope and vanish on $\partial M$. These are in fact the images of the forms $\phi_{ij}^{\prime}\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$ under the map $\Psi_{f}$, where $\phi_{ij}^{\prime}:M\to\mathfrak{gl}(n)$ has the same continuity and smoothness as the maps $\phi^{k}_{lij}$.
@@
Note that if $\hat{\phi}\in\mathcal{A}(f)$, then $\hat{\phi}\in\mathcal{A}(f\cdot h)$ for any continuous map $h:M\to GL(n)$, so the dependence of $\mathcal{A}(f)$ on the choice of frame is limited to dependence on its discontinuities. Additionally, the definition of compatible test forms does not depend on the choice of tangent frames $E_{e}$ and $E_{p}$, because if $\tilde{E}_{e}$ is another choice of orthonormal tangent frame for $e$, then $\tilde{E}_{e}^{T}=E_{e}^{T}\cdot\begin{bmatrix}h&0\\ 0&1\end{bmatrix}$ for some orthogonal matrix $h\in O(n-1)$. Therefore, we have on $\Phi^{-1}(\mathring{e})$,
@@
$[\hat{\phi}^{T}]_{\tilde{E}_{e}^{T}}$ $=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\right)\left([\hat{\phi}^{T}]_{E_{e}^{T}}\right)=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([\hat{\phi}^{T^{\prime}}]_{E_{e}^{T^{\prime}}}\right)$
@@
$=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\begin{bmatrix}h&0\\ 0&1\end{bmatrix}\right)\left([\hat{\phi}^{T^{\prime}}]_{\tilde{E}_{e}^{T^{\prime}}}\right)$
@@
$=\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([\hat{\phi}^{T^{\prime}}]_{\tilde{E}_{e}^{T^{\prime}}}\right).$
@@
Similar reasoning shows that if an alternative tangential orthonormal frame $\tilde{E}_{p}$ is chosen for $p$, then equation (20) is valid for $\hat{\phi}^{T}]_{\tilde{E}_{p,e}^{T}}$ as well.
[…]
In summary, there is a bijection $\Psi_{f}:\mathcal{C}(\mathcal{T},M)\to\mathcal{A}(f)$ for any compatible frame $f$ which is induced by the pointwise isometry $\Psi_{y}^{T}:\mathfrak{gl}(n)\to\mathrm{End}(T_{\Phi^{T}(y)}M)$.
@@
With very similar reasoning, there exist bijections $\Psi_{E_{e}^{T}}:C^{0}\Omega^{k}(e;\mathfrak{gl}(n))\to C^{0}\Omega^{k}(e;\mathrm{End}(TM))$ and $\Psi_{E_{p,e}^{T}}:C^{0}\Omega^{k}(p;\mathfrak{gl}(n))\to C^{0}\Omega^{k}(p;\mathrm{End}(TM))$ defined in exactly the same way, but using the frames $E_{e}^{T}$ and $E_{p,e}^{T}$ defined on $e$ and $p$ respectively.
@@
We can define the following endomorphism-valued forms, which are a $2$-form defined on the codimension-0 polytope $T$, a pair of $1$-forms defined on the codimension-1 polytope $e=T\cap T^{\prime}$, and a 0-form defined on the codimension-2 polytope $p\subset T\cap e$ respectively:
@@
$\hat{\Theta}_{p}^{T}$ $:=\Psi_{E_{p,e}^{T}}(\Theta_{p}w^{n-1}_{n})=\Theta_{p}(\vec{n}\otimes\vec{\nu}^{\flat}-\vec{\nu}\otimes\vec{n}^{\flat}).$
@@
On each $T\subset M$, $\hat{R}$ is the usual Riemann curvature tensor, i.e. it is the $\operatorname{End}(TM)$-valued 2-form given by
@@
On each $e\subset T$, $\hat{\mathrm{I\!I}}_{e}^{T}$ is defined as follows. For all $x\in\mathring{e}$, $X,Y\in T_{x}M$, and $Z\in T_{x}e$,
@@
The formulas for $\hat{R}$ and $\hat{\Theta}_{p}^{T}$ are clear from their definitions. To derive the formula for $\hat{\mathrm{I\!I}}_{e}^{T}$, let $E_{i}$ be shorthand for $(E_{e}^{T})_{i}$, and let $X=X^{i}E_{i}$, $Y=Y^{i}E_{i}$, and $Z=Z^{i}E_{i}$. Assume that $Z$ is tangent to $e$, so that $Z^{n}=0$. Then
@@
If $f$ is any compatible frame and $\hat{\phi}\in\mathcal{A}(f)$, then we define
@@
The fact that $\hat{R}_{\mathrm{dist}}$ is well-defined is a consequence of the following theorem.
@@
If $\hat{\phi}\in\mathcal{A}(f)$ for some compatible frame $f$, then
$\left\langle\left\langle\hat{R}_{\mathrm{dist}},\hat{\phi}\right\rangle\right\rangle=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle\hat{R}\wedge\hat{\phi}\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left[\!\left[\left\langle\hat{\mathrm{I\!I}}_{e}\wedge i_{\mathring{e}}^{*}\hat{\phi}\right\rangle\right]\!\right]+\sum_{p\subset M}\int_{\mathring{p}}\left\langle\hat{\Theta}_{p}^{T}\wedge i_{\mathring{p}}^{*}\hat{\phi}^{T}\right\rangle.$ (22)
Here $\left[\!\left[\left\langle\hat{\mathrm{I\!I}}_{e}\wedge i_{\mathring{e}}^{*}\phi\right\rangle\right]\!\right]$ is defined as $\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\phi^{T}\right\rangle-\left\langle\hat{\mathrm{I\!I}}_{e}^{T^{\prime}}\wedge i_{\mathring{e}}^{*}\hat{\phi}^{T^{\prime}}\right\rangle$, which is interpreted as the jump in second fundamental form, while $\left\langle\Theta_{p}\wedge i_{\mathring{p}}^{*}\phi\right\rangle$ is defined as $\left\langle\Theta_{p}^{T}\wedge i_{\mathring{p}}^{*}\phi^{T}\right\rangle$, where $T$ is any of the $n$-dimensional polytopes containing $p$ as a face.

### v2

@@ 2.5 Removing Frame-Dependency
While useful for analysis, (4) is inconvenient for many applications because it requires knowledge of a specific compatible frame, which are not easy to construct or represent. The purpose of this subsection is to remove this dependency by expressing the distributional curvature in terms of endomorphism-valued forms. The cost is that the test forms must be discontinuous and metric-dependent.
First, we can define a symmetric nondegenerate bilinear form on the vector space $\mathrm{End}(T_{x}M)=T_{x}M\otimes T^{*}_{x}M$ by setting
$\left\langle\alpha\otimes v,\beta\otimes w\right\rangle:=\alpha(w)\beta(v)$ (17)
for $\alpha,\beta\in T_{x}^{*}(M)$, $v,w\in T_{x}M$ and extending multilinearly. Expressed in a basis $\{F_{i}\}_{i=1}^{n}$ and its dual basis $\{F^{i}\}_{i=1}^{n}$, this would be $\left\langle A^{i}_{j}F_{i}\otimes F^{j},B^{k}_{l}F_{k}\otimes F^{l}\right\rangle=A^{i}_{j}B^{j}_{i}$. Therefore, picking any basis for $T_{x}M$ gives a linear isometry between $\mathfrak{gl}(n)$ and $\mathrm{End}(T_{x}M)$, by setting $F^{j}(\hat{\phi}^{T}(F_{i})):=\phi^{j}_{i}$. The left-hand side of this equation could be more succinctly written $([\hat{\phi}]_{F})^{j}_{i}$, where $\hat{\phi}]_{F}$ is the matrix representation of the linear map $\hat{\phi}:T_{x}M\to T_{x}M$ in the basis $F$. In particular, for $y\in B_{T}$, the orthonormal basis $F^{T}(y)$ gives a special isometry $\Psi_{y}^{T}:\mathfrak{gl}(n)\to\mathrm{End}(T_{\Phi^{T}(y)}M)$.
This pointwise isometry in turn defines a map $\Psi_{f}:\bigsqcup_{T\subseteq M}C^{0}\Omega^{n-2}(\mathring{T};\mathfrak{gl}(n))\to\bigsqcup_{T\subseteq M}C^{0}\Omega^{n-2}(\mathring{T};\mathrm{End}(TM))$, by setting $\Psi_{f}(\phi)|_{x}$ equal to $\Psi_{{\Phi^{T}}^{-1}(x)}^{T}(\phi|_{x})$ if $x$ is in $\mathring{T}$. Clearly, because $\Phi_{y}^{T}$ is invertible for all $y\in\mathring{B}_{T}$ and $\Phi^{T}$ is a diffeomorphism when restricted to $\mathring{B}_{T}$, $\Psi_{f}$ is a bijection, with inverse given by $\Psi_{f}^{-1}(\hat{\phi})|_{x}:=(\Psi_{{\Phi^{T}}^{-1}(x)}^{T})^{-1}(\hat{\phi}|_{x})$.
Definition 6.
Let $f$ be a compatible frame. A compatible $\mathrm{End}(TM)$-valued test form is the image of a compatible $\mathfrak{so}(n)$-valued test form (see Definition 5) under the map $\Psi_{f}$. The space of compatible $\mathrm{End}(TM)$-valued test forms is denoted $\mathcal{A}(f,\mathcal{T},M):=\Psi_{f}(\mathcal{C}(\mathcal{T},M))$.
Several properties of $\mathrm{End}(TM)$-valued test forms make them more suited to frame-independent computations.
@@
Since $F^{T}$ is continuous and Lipschitz on $B_{T}$ and ${\Phi^{T}}^{*}\phi|_{\mathring{T}}$ has a unique Lipschitz extension $\tilde{\phi}^{T}$ on $B_{T}$, there exists a unique Lipschitz extension of ${\Phi^{T}}^{*}\hat{\phi}|_{\mathring{T}}$ to $B_{T}$, which at each point $y\in B_{T}$ is equal to $\Psi_{y}^{T}(\tilde{\phi}^{T}|_{y})$. We will use $\hat{\phi}^{T}$ to refer to this extension of ${\Phi^{T}}^{*}\hat{\phi}|_{\mathring{T}}$. Similarly to the case for compatible $\mathfrak{so}(n)$-valued test forms, the pullback $i_{\mathring{e}}^{*}\hat{\phi}|_{T}:={\Phi^{T}}^{-*}i_{{\Phi^{T}}^{-1}(\mathring{e})}^{*}\hat{\phi}^{T}$ is a well defined continuous extension of $\hat{\phi}|_{\mathring{T}}$.
The two key properties we will need about compatible $\mathrm{End}(TM)$-valued test forms are proved in the following lemma.
Lemma 7.
Let $\hat{\phi}=\Psi_{f}(\phi)$ be a compatible $\mathrm{End}(TM)$-valued test form. Then the following change of basis equation holds for each $T,T^{\prime}\subseteq M$ and $e=T\cap T^{\prime}$:
$i_{\mathring{e}}^{*}\hat{\phi}|_{T}]_{E_{e}^{T}}=\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}]_{E_{e}^{T^{\prime}}}\right).$ (19)
Additionally, the following equation holds for each $p\subset e\subset T$ such that $E_{p,e}^{T}$ has an inward-pointing normal vector as its lasty entry:
${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}\left(\exp\left(s~r_{p,e}^{T}w^{n-1}_{n}\right)\right)\left([{\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}\right).$ (20)
Here, $\chi_{s,p}^{T}$ is the map defined within condition 3 of Definition 5.
Proof.
By (18), at points $x\in e=T\cap T^{\prime}$, we have on $T$ that $\mathrm{Ad}((\mu_{e}^{T})^{-1})([i_{\mathring{e}}^{*}\hat{\phi}|_{T}]_{E_{e}^{T}})=i_{\mathring{e}}^{*}\phi=\mathrm{Ad}((\mu_{e}^{T^{\prime}})^{-1})([i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}]_{E^{T^{\prime}}_{e}})$, since $\phi$ is compatible. However, because $\mu_{e}^{T^{\prime}}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\mu_{e}^{T}$, we have
$\mathrm{Ad}((\mu_{e}^{T})^{-1})\left([i_{\mathring{e}}^{*}\hat{\phi}|_{T}]_{E_{e}^{T}}-\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}|_{T^{\prime}}]_{E^{T^{\prime}}_{e}}\right)\right)=0$
By applying $\mathrm{Ad}(\mu_{e}^{T})$ to both sides of this equation, we get Equation (19).
Similarly, the form $i_{\mathring{p}}^{*}\hat{\phi}^{T}]_{F^{T}}=i_{\mathring{p}}^{*}\phi$ is well-defined, in the sense that ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}$ does not depend on $s$ or $T$. Below we will use the shorthand $F(s,x)=F^{T}\circ\psi_{p}^{T}(s,x)$.
@@
${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}\left(\exp\left(s~r_{p,e}^{T}w^{n-1}_{n}\right)A_{p,e}^{T}\mu_{e}^{T}\right)\left([{\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{F(s,\cdot)}\right)$
@@
Since ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}={\chi_{0,p}^{T}}^{*}\tilde{\phi}^{T}$ is one of the defining characteristics of a compatible $\mathfrak{so}(n)$-valued test form and ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{F(s,\cdot)}={\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}$, the second equation can be substituted into the first to obtain Equation (20).
A more intuitive way to understand what this lemma gives us is: rather than $\hat{\phi}$ itself being continuous, we get that the matrix representations of $\hat{\phi}$ in special frames adapted to the mesh are only allowed to have discontinuties in a special way, and except for the dependence on $r_{p,e}^{T}$, the discontinuities do not depend on the frame $f$.
@@
Just as for compatible test forms, an important class of compatible $\mathrm{End}(TM)$-valued test forms is given by the forms $\phi^{k}_{lij}(f_{k}\otimes f^{*}\theta^{l})\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$, where the coefficients $\phi^{k}_{lij}$ are continuous compactly supported functions on $M$ which are $C^{2}$ on the interior of each codimension-0 polytope and vanish on $\partial M$ and which alternate in the $k,l$ indices. These are exactly the images of the forms $\phi_{ij}^{\prime}\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$ under the map $\Psi_{f}$, where $\phi_{ij}^{\prime}:M\to\mathfrak{so}(n)$ has the same continuity and smoothness as the maps $\phi^{k}_{lij}$. Unlike the case for compatible $\mathfrak{so}(n)$-valued test forms, the set of compatible $\mathrm{End}(TM)$-valued test forms is in general not a superset of the smooth $\mathrm{End}(TM)$-valued test forms, as discontinuities in $f$ can force discontinuties in $\hat{\phi}$.
@@
Note that if $\hat{\phi}\in\mathcal{A}(f,\mathcal{T},M)$, then $\hat{\phi}\in\mathcal{A}(f\cdot h,\mathcal{T},M)$ for any continuous piecewise-$C^{2}$ map $h:M\to O(n)$, so the dependence of $\mathcal{A}(f,\mathcal{T},M)$ on the choice of frame is limited to dependence on its discontinuities. Additionally, the validity of equations (19) and (20) does not depend on the choice of tangent frames $E_{e}$ and $E_{p}$, because if $\tilde{E}_{e}$ is another choice of orthonormal tangent frame for $e$, then $\tilde{E}_{e}^{T}=E_{e}^{T}\cdot\begin{bmatrix}h&0\\ 0&1\end{bmatrix}$ for some orthogonal matrix $h\in O(n-1)$. Therefore, we have on $\Phi^{-1}(\mathring{e})$,
@@
$[i_{\mathring{e}}^{*}\hat{\phi}^{T}]_{\tilde{E}_{e}^{T}}$ $=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}^{T}]_{E_{e}^{T}}\right)=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}^{T^{\prime}}]_{E_{e}^{T^{\prime}}}\right)$
@@
$=\mathrm{Ad}\left(\begin{bmatrix}h^{-1}&0\\ 0&1\end{bmatrix}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\begin{bmatrix}h&0\\ 0&1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}^{T^{\prime}}]_{\tilde{E}_{e}^{T^{\prime}}}\right)$
@@
$=\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)\left([i_{\mathring{e}}^{*}\hat{\phi}^{T^{\prime}}]_{\tilde{E}_{e}^{T^{\prime}}}\right).$
@@
Similar reasoning shows that if an alternative (similarly oriented) tangential orthonormal frame $\tilde{E}_{p}$ is chosen for $p$, then equation (20) is valid for ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{\tilde{E}_{p,e}^{T}}$ as well.
@@
We may also define bijections $\Psi_{E_{e}^{T}}:C^{0}\Omega^{k}(e;\mathfrak{gl}(n))\to C^{0}\Omega^{k}(e;\mathrm{End}(TM))$ and $\Psi_{E_{p,e}^{T}}:C^{0}\Omega^{k}(p;\mathfrak{gl}(n))\to C^{0}\Omega^{k}(p;\mathrm{End}(TM))$ in exactly the same way as $\Psi_{f}$, but using the frames $E_{e}^{T}$ and $E_{p,e}^{T}$ defined on $e$ and $p$ respectively.
@@
We can then define the following endomorphism-valued forms, which are a $2$-form defined on the codimension-0 polytope $T$, a pair of $1$-forms defined on the codimension-1 polytope $e=T\cap T^{\prime}$, and a 0-form defined on the codimension-2 polytope $p\subset T\cap e$ respectively:
@@
$\hat{\Theta}_{p}^{T}$ $:=\Psi_{E_{p,e}^{T}}(\Theta_{p}w^{n-1}_{n})=\Theta_{p}(-\vec{n}\otimes\vec{\nu}^{\flat}+\vec{\nu}\otimes\vec{n}^{\flat}).$
Note that for consistency, when defining $\hat{\Theta}_{p}^{T}$, we must take care to use the face $e$ such that $E_{p,e}^{T}$ has $\vec{n}^{T}$ as its last vector, so that the rotation from $E_{p,e}^{T}$ to $E_{p,e^{\prime}}^{T}$ ($e^{\prime}$ being the other codimension-1 face meeting $p$) is in the counterclockwise direction. Otherwise, as mentioned at the end of section 2.3, the sign of $\Theta_{p}$ will need to flip.
@@
Here $\left\langle\cdot,\cdot\right\rangle_{T}$ and $\nabla^{T}$ refer to the inner product and covariant derivative induced by the metric $g^{T}$, respectively.
@@
On each $T\subseteq M$, $\hat{R}$ is the usual Riemann curvature tensor, i.e. it is the $\operatorname{End}(TM)$-valued 2-form given by
@@
On each $e\subset T$, $\hat{\mathrm{I\!I}}_{e}^{T}$ is given as follows. For all $x\in\mathring{e}$, $X,Y\in T_{x}M$, and $Z\in T_{x}e$,
@@
The claims about $\hat{R}$ and $\hat{\Theta}_{p}^{T}$ are clear from their definitions. To derive the formula for $\hat{\mathrm{I\!I}}_{e}^{T}$, let $E_{i}$ be shorthand for $(E_{e}^{T})_{i}$, and let $X=X^{i}E_{i}$, $Y=Y^{i}E_{i}$, and $Z=Z^{i}E_{i}$. Assume that $Z$ is tangent to $e$, so that $Z^{n}=0$. Then
@@
If $f$ is any compatible frame and $\hat{\phi}\in\mathcal{A}(f,\mathcal{T},M)$, then we define
@@
The following theorem gives a more explicit formula for $\hat{R}_{\mathrm{dist}}$ and demonstrates that $\hat{R}_{\mathrm{dist}}$ is well-defined.
@@
If $\hat{\phi}\in\mathcal{A}(f,\mathcal{T},M)$ for some compatible frame $f$, then
$\left\langle\left\langle\hat{R}_{\mathrm{dist}},\hat{\phi}\right\rangle\right\rangle=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle\hat{R}\wedge\hat{\phi}\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left[\!\left[\left\langle\hat{\mathrm{I\!I}}_{e}\wedge i_{\mathring{e}}^{*}\hat{\phi}\right\rangle\right]\!\right]+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\left\langle\hat{\Theta}_{p}\wedge i_{\mathring{p}}^{*}\hat{\phi}\right\rangle$. (22)
[…]